# An efficient probabilistic hardware architecture for diffusion-like models

### v1

@@ D.3.2 Initialization and readout
D.4 Analysis of a complete sampling program
@@
This exact procedure was used to estimate the energy consumption of the MEBMs in Fig. 1 in the article. In this case, $T=1$ and $K$ were estimated from the autocorrelation data for each layer; see section  K.
D.5 Level of realism
@@
Some of the discrepancies between the high-level and detailed model can be attributed to overheads associated with real circuits. A real implementation of the biasing circuit discussed in section  D.D.1 is more complicated than the theoretical model because tunable resistors do not exist. Communications with neighboring cells over long wires require driver circuits, which consume additional energy beyond what is spent charging the line. Despite this, real circuits are bound by the same fundamental physics as the simplified models presented here. As such, the simplified models tend to estimate energy consumption within a factor of two or three of real-life values.
@@
D.5.1 Programming the weights and biases
Section D.D.1 discusses a simple circuit that uses resistors to implement the multiply-accumulate required by the conditional update rule. Key to this is being able to tune the conductance of the resistors to implement specific sets of weights and biases (see Eq. (86)).
@@ D.5.1 Programming the weights and biases
D.5.2 Off-chip communication
External devices have to communicate with our chip for it to be useful. The cost of this communication depends strongly on the tightness of integration between the two systems and is impossible to reason about at an abstract level. As such, the analysis of communication here (as in Section D.3.2) was limited to the cost of getting bits out to the edge of our chip, which is a lower bound on the actual cost.
However, we have found that a more detailed analysis, which includes the cost of communication between two chips mediated by a PCB, does not significantly change the results at the system level. The fundamental reason for this is that sampling programs for complex models run for many iterations before mixing and sending the results back to the outside world. This is reflected in the discrepancy between $E_{\text{samp}}$ and $E_{\text{init}}+E_{\text{read}}$ found in section D.D.4.
D.5.3 Supporting circuitry
@@
Appendix E Energy analysis of GPUs
@@
The models were derived from available implementations and are based on ResNet [3] and UNet [11] style architectures. Their FID performance is consistent with published literature values [2, 1, 9]. The goal is not to achieve state of the art performance, but to represent the relative scales of energy consumption of the algorithms.
@@
Appendix F Total correlation penalty
@@
F.1 Gradients of the total correlation penalty
@@
F.2 Low dimensional embedding of the data
For all experiments in this article, the embedding function $f$ used in autocorrelation calculations was the encoding neural network used in FID computation. Using the encoder was an arbitrary choice, and we could have just as easily used a much simpler function. For example, we found that random linear projections $y[j]=Ax[j]$ worked just as well as the neural network for autocorrelation calculations.
F.3 Control of the penalty strength
@@
To address this, we employ an Adaptive Correlation Penalty (ACP) scheme that dynamically adjusts $\lambda_{t}$ based on an estimate of the model’s current mixing time. We use the autocorrelation of the Gibbs sampling chain, $r_{yy}^{t}$, as a proxy for mixing, as described in Section H and the main text, Eq. 18.
@@
Figure 14: The active correlation penalty The dynamics of $r_{yy}^{t}$ and $\lambda_{t}$ over a training run. Large values of $r_{yy}^{t}$ lead to increasingly large values of $\lambda_{t}$, which cause $r_{yy}^{t}$ to decrease. The system reaches a stable configuration by the end of training.
Appendix G Embedding integers into Boltzmann machines
@@
Appendix H The autocorrelation function and mixing time
This section gives a brief derivation of how the Mixing time of a Markov chain can be estimated using its autocorrelation. Proofs of the properties noted here can be found in most standard textbooks on Markov chains, such as [6].
Suppose $X$ is a discrete-time Markov chain on a finite state space $\{1,\ldots,d\}$. Suppose its transition kernel is time-homogeneous and given by a matrix $P=\left(p_{xy}\right)_{1\leq x,y\leq d}$ with entries $p_{xy}=\mathbb{P}\left(X_{t+1}=y|X_{t}=x\right)$. Furthermore, assume the Markov chain is:
  • irreducible, i.e. it is possible to get from any starting node to any other node in a finite number of steps, and
  • aperiodic, i.e. for any starting position $x\in\{1,\ldots,d\}$, there exists some $T\in\mathbb{N}$ such that for all $t\geq T$, $\mathbb{P}\left(X_{t}=x|X_{0}=x\right)>0$.
Since this Markov chain is defined on a finite state space and is irreducible, there exists a unique stationary distribution $\pi=\pi P$. Furthermore, as a consequence of aperiodicity, for any starting distribution $\psi_{0}$, the distribution $\psi_{t}$ of the Markov chain at time $t$ converges to $\pi$ as $t\rightarrow\infty$, that is
$\psi_{t}=\psi_{0}P^{t}\rightarrow\pi\;\text{ as }\;t\rightarrow\infty.$
The transition matrix can be diagonalized as $P=U^{-1}\Sigma\,U$, where $\Sigma$ is a diagonal matrix and without loss of generality, we can assume that its entries are ordered $1=\sigma_{1}>\sigma_{2}\geq\ldots\geq\sigma_{d}\geq 0$. Then the $t$th power of $P$ can be written as $P^{t}=U^{-1}\Sigma^{t}U$. We write $U(i,x)$ for the entry in the $i$th row and $x$th column of $U$ (and likewise for $U^{-1}$). The first left eigenvector of $P$ (and hence the first row of $U$) is the stationary distribution $\pi=\pi P=U(1,\cdot)$. The first right eigenvector of $P$ is a column of ones, that is $U^{-1}(\cdot,1)=\mathbf{1}$.
Let $f:\{1,\ldots,d\}\rightarrow\mathbb{R}$ be any function and write,
$\mu^{f}_{0}\equiv\mathbb{E}_{Y\sim\psi_{0}}\left[f(Y)\right]=\mathbb{E}\left[f(X_{0})\right]$ (104)
where $\psi_{0}$ is the initial distribution of the Markov chain $X$ and,
$\mu^{f}_{\infty}\equiv\mathbb{E}_{Y\sim\pi}\left[f(Y)\right]=\lim_{t\rightarrow\infty}\mathbb{E}\left[f(X_{t})\right]$ (105)
Then, write $\delta_{k}$ for the column vector with a one at $k$th entry and zeros elsewhere. We can then compute,
$\begin{split}\mathbb{E}\left[f(X_{0})f(X_{t})\right]&=\sum_{x_{0}=1}^{d}f(x_{0})\mathbb{P}\left[X_{0}=x_{0}\right]\mathbb{E}\left[f(X_{t})|X_{0}=x_{0}\right]=\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,(\delta^{T}_{x_{0}}P^{t})(x)\\[5.69054pt] &=\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\psi_{0}(x_{0})\,\sum_{j=1}^{d}U^{-1}(x_{0},j)\,\sigma_{j}^{t}\,U(j,x)\\[5.69054pt] &=\left(\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\psi_{0}(x_{0})U^{-1}(x_{0},1)\,\sigma_{1}^{t}\,U(1,x)\right)+\\ &\quad+\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\psi_{0}(x_{0})\,\sum_{j=2}^{d}U^{-1}(x_{0},j)\,\sigma_{j}^{t}\,U(j,x)\\[5.69054pt] &=\left(\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\psi_{0}(x_{0})\pi(x)\right)+\sum_{j=2}^{d}\sigma_{j}^{t}\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\psi_{0}(x_{0})\,U^{-1}(x_{0},j)\,U(j,x)\\[5.69054pt] &=\mu^{f}_{0}\mu^{f}_{\infty}+\sum_{j=2}^{d}\sigma_{j}^{t}c_{j},\end{split}$ (106)
where $c_{j}$ are constants independent of $t$. Hence, if we can plot the quantity $\operatorname*{\mathbb{E}}{f(X_{0})f(X_{t})}-\mu^{f}_{0}\mu^{f}_{\infty}$ for very large $t$, the contributions of the smaller eigenvalues become negligible relative to the contribution of $\sigma_{2}$, allowing us to estimate the value of $\sigma_{2}$.
Why does knowing $\sigma_{2}$ help us compute the mixing time? Recall that the mixing time $\tau$ is defined as
$\|\mu-\nu\|_{\text{TV}}=\frac{1}{2}\sum_{y\in\mathcal{X}}|\mu(y)-\nu(y)|$ denotes the total variation distance, and $\varepsilon>0$ is a prescribed tolerance.
We can rewrite the total variation distance in this definition as
[…]
$=\frac{1}{2}\sum_{x=1}^{d}\left|\pi(x)-\pi(x)+\sum_{j=2}^{d}\big(\psi_{0}\cdot U^{-1}(\cdot,j)\big)\,\sigma_{j}^{t}\,U(j,x)\right|$
$\leq\frac{1}{2}\sum_{x=1}^{d}\sum_{j=2}^{d}\big|\psi_{0}\cdot U^{-1}(\cdot,j)\big|\,|\sigma_{j}^{t}|\,\big|U(j,x)\big|$
$=\sum_{j=2}^{d}\sigma_{j}^{t}\,a_{j}\leq\sigma^{t}_{2}\sum_{j=2}^{d}a_{j}$
where $a_{j}$ are some non-negative constants. Therefore, we can establish the following upper bound:
The smaller the value of $\sigma_{2}$, the faster the mixing time.
Appendix I Deterministic embeddings for DTMs
@@
In Section V of the paper, we mention hybrid thermodynamic models, the purpose of which is to combine the flexibility of classical neural networks (NNs) with the efficiency of probabilistic computers. For example, in the context of image generation, a small convolutional neural network can be used to map color images into a format compatible with a binary DTM. To properly take advantage of the DTM’s energy efficiency, the classical model should be at least 2 or 3 orders of magnitude smaller (e.g., in terms of parameter count or number of operations per sample) than the DTM.
@@
There are various options for the type of classical model one can use for the embedding, e.g., invertible models such as GLOW [5] or Normalizing Flows [10], as well as simpler solutions, such as an Autoencoder.
@@
Our hybrid model achieved a FID score of $\sim 60$ on CIFAR10. Our DTM had 8 million parameters, the decoder had 65k, and the encoder and critic were both below 500k parameters. At inference time, only the DTM and the decoder are used. To achieve a similar performance with a classical GAN, the decoder/generator requires about 500000 parameters.
Appendix J Some details on our RNG
@@
Appendix K MEBM experiments
Our experiments on MEBMs were conducted in the typical way [12]. We employed the same Boltzmann machine architecture as we typically use for the DTM layers, specifically $L=70$ with $G_{12}$ connectivity. Random nodes were chosen to represent the data, and the rest were left as latent variables (as discussed in section C).
@@
The exponential decay rates extracted from Fig. 16 were used as the mixing times in Fig. 2 in the article. Calling this a "mixing time" is a slight abuse of nomenclature. However, we did not think it made enough of a difference to the article’s message to disambiguate (since it is an upper bound on the mixing time, as discussed in Section H).

### v2

@@ E.3.2 Initialization and readout
E.4 Analysis of a complete sampling program
@@
This exact procedure was used to estimate the energy consumption of the MEBMs in Fig. 1 in the article. In this case, $T=1$ and $K$ were estimated from the autocorrelation data for each layer; see section  L.
E.5 Level of realism
@@
Some of the discrepancies between the high-level and detailed model can be attributed to overheads associated with real circuits. A real implementation of the biasing circuit discussed in section  E.E.1 is more complicated than the theoretical model because tunable resistors do not exist. Communications with neighboring cells over long wires require driver circuits, which consume additional energy beyond what is spent charging the line. Despite this, real circuits are bound by the same fundamental physics as the simplified models presented here. As such, the simplified models tend to estimate energy consumption within a factor of two or three of real-life values.
@@
E.5.1 Programming the weights and biases
Section E.E.1 discusses a simple circuit that uses resistors to implement the multiply-accumulate required by the conditional update rule. Key to this is being able to tune the conductance of the resistors to implement specific sets of weights and biases (see Eq. (83)).
@@ E.5.1 Programming the weights and biases
E.5.2 Off-chip communication
External devices have to communicate with our chip for it to be useful. The cost of this communication depends strongly on the tightness of integration between the two systems and is impossible to reason about at an abstract level. As such, the analysis of communication here (as in Section E.3.2) was limited to the cost of getting bits out to the edge of our chip, which is a lower bound on the actual cost.
However, we have found that a more detailed analysis, which includes the cost of communication between two chips mediated by a PCB, does not significantly change the results at the system level. The fundamental reason for this is that sampling programs for complex models run for many iterations before mixing and sending the results back to the outside world. This is reflected in the discrepancy between $E_{\text{samp}}$ and $E_{\text{init}}+E_{\text{read}}$ found in section E.E.4.
E.5.3 Supporting circuitry
@@
Appendix F Energy analysis of GPUs
@@
The models were derived from available implementations and are based on small versions of ResNet [3] and UNet [11] style architectures. Their FID performance is consistent with published literature values [2, 1, 9]. The goal is not to achieve state of the art performance, but to represent the relative scales of energy consumption of the algorithms. For a direct simulation of the Ising models on a GPU, theoretical efficiency on the order of $10^{-4}$ joules per sample comparable in performance and efficiency to the small VAE.
@@
Appendix G Autocorrelation and mixing time
We monitor the mixing of our Gibbs samplers via the autocorrelation of a low-dimensional projection of the Markov chain state.
Let $\{x[j]\}_{j\geq 0}$ be a discrete-time Markov chain on a finite state space $\{1,\ldots,d\}$, with time-homogeneous transition kernel $P=(p_{xy})_{1\leq x,y\leq d}$ given by
$p_{xy}=\mathbb{P}(x[j+1]=y\mid x[j]=x).$
In our setting, $x[j]$ is the state of the Boltzmann machine at Gibbs iteration $j$.
We fix a function (projection) $f:\{1,\ldots,d\}\rightarrow\mathbb{R}$ and define the scalar observable
$y[j]=f(x[j]).$
We then define the mean
$\mu\equiv\mathbb{E}\bigl[y[j]\bigr]$ (94)
and the (normalized) autocorrelation function
$r_{yy}[k]=\frac{\mathbb{E}\bigl[(y[j]-\mu)\,(y[j+k]-\mu)\bigr]}{\mathbb{E}\bigl[(y[j]-\mu)^{2}\bigr]},\qquad k\in\mathbb{N}.$ (95)
Here $\mathbb{E}[\cdot]$ denotes expectation with respect to the joint law of the Markov chain. In practice, we approximate this expectation by averaging over time and across multiple independent Gibbs sampling chains.
For all experiments in this article, we take $f$ to be the encoder network used in the FID computation, applied to the visible states of the Boltzmann machine. This choice is largely arbitrary: we found that much simpler embeddings, such as random linear projections $y[j]=Ax[j]$, behave similarly well for the purposes of autocorrelation-based mixing diagnostics.
Spectral decomposition and decay of autocorrelation
We now briefly recall how the decay rate of autocorrelation is tied to the spectral properties of the transition kernel $P$, and hence to the mixing time of the Markov chain. Proofs of the statements below can be found in standard references such as [6].
Assume the Markov chain $\{x[j]\}_{j\geq 0}$ is
  • irreducible, i.e. any state can be reached from any other in a finite number of steps, and
  • aperiodic, i.e. for any $x\in\{1,\ldots,d\}$ there exists $T\in\mathbb{N}$ such that for all $t\geq T$, $\mathbb{P}(x[t]=x\mid x[0]=x)>0$.
Since the chain is finite and irreducible, there exists a unique stationary distribution $\pi$ satisfying
$\pi=\pi P.$
Furthermore, aperiodicity implies that, for any initial distribution $\psi_{0}$ of $x[0]$, the law $\psi_{t}$ of $x[t]$ converges to $\pi$ as $t\to\infty$:
$\psi_{t}=\psi_{0}P^{t}\;\longrightarrow\;\pi\qquad\text{as }t\rightarrow\infty.$
Assume further that $P$ is diagonalizable with real eigenvalues (which is always true if the chain is reversible, like in the case of Gibbs sampling), and write its eigendecomposition as
$P=U^{-1}\Sigma\,U,$
where $\Sigma$ is diagonal, and its entries are ordered
$1=\sigma_{1}>\sigma_{2}\geq\cdots\geq\sigma_{d}\geq 0.$
We denote by $U(i,x)$ the entry in the $i$th row and $x$th column of $U$ (and similarly for $U^{-1}$). The first left eigenvector of $P$ (the first row of $U$) is the stationary distribution,
$\pi=\pi P=U(1,\cdot),$
and the first right eigenvector is a column of ones,
$U^{-1}(\cdot,1)=\mathbf{1}.$
Let $f:\{1,\ldots,d\}\to\mathbb{R}$ be the same function as above, and write
$\mu^{f}_{0}\equiv\mathbb{E}_{Y\sim\psi_{0}}\bigl[f(Y)\bigr]=\mathbb{E}\bigl[f(x[0])\bigr],\qquad\mu^{f}_{\infty}\equiv\mathbb{E}_{Y\sim\pi}\bigl[f(Y)\bigr]=\lim_{t\rightarrow\infty}\mathbb{E}\bigl[f(x[t])\bigr],$ (96)
where $\psi_{0}$ is the initial distribution of $x[0]$.
Let $\delta_{k}$ denote the column vector with a 11 in position $k$ and 0 elsewhere. Using the eigendecomposition of $P$, the transition probabilities can be written as
$(\delta_{x_{0}}^{T}P^{t})(x)=\sum_{j=1}^{d}U^{-1}(x_{0},j)\,\sigma_{j}^{t}\,U(j,x).$
Thus,
$\begin{split}\mathbb{E}\bigl[f(x[0])f(x[t])\bigr]&=\sum_{x_{0}=1}^{d}f(x_{0})\,\mathbb{P}\bigl[x[0]=x_{0}\bigr]\,\mathbb{E}\bigl[f(x[t])\mid x[0]=x_{0}\bigr]\\ &=\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,(\delta_{x_{0}}^{T}P^{t})(x)\\ &=\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,\sum_{j=1}^{d}U^{-1}(x_{0},j)\,\sigma_{j}^{t}\,U(j,x)\\ &=\left(\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,U^{-1}(x_{0},1)\,\sigma_{1}^{t}\,U(1,x)\right)\\ &\quad+\sum_{j=2}^{d}\sigma_{j}^{t}\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,U^{-1}(x_{0},j)\,U(j,x).\end{split}$ (97)
Using $U^{-1}(\cdot,1)=\mathbf{1}$ and $U(1,\cdot)=\pi$, the first term simplifies to
$\sum_{x_{0}=1}^{d}\sum_{x=1}^{d}f(x_{0})f(x)\,\psi_{0}(x_{0})\,\pi(x)=\mu^{f}_{0}\mu^{f}_{\infty}.$
We can therefore write
$\mathbb{E}\bigl[f(x[0])f(x[t])\bigr]=\mu^{f}_{0}\mu^{f}_{\infty}+\sum_{j=2}^{d}\sigma_{j}^{t}c_{j},$ (98)
where the coefficients $c_{j}$ are constants (depending on $f$ and $\psi_{0}$ but not on $t$). For large $t$, the terms with smaller eigenvalues become negligible compared to the contribution from $\sigma_{2}$, so the covariance
$\mathbb{E}\bigl[f(x[0])f(x[t])\bigr]-\mu^{f}_{0}\mu^{f}_{\infty}$
decays asymptotically like $\sigma_{2}^{t}$. In particular, under stationarity ($\psi_{0}=\pi$ so that $\mu^{f}_{0}=\mu^{f}_{\infty}=\mu$), the autocorrelation $r_{yy}[t]$ defined in (95) decays approximately as
$r_{yy}[t]\approx C\,\sigma_{2}^{t}$
for some constant $C$ depending on $f$.
From eigenvalues to mixing time
The spectral quantity $\sigma_{2}$ is directly related to the mixing time of the Markov chain. Recall that the mixing time is defined by
where
$\|\mu-\nu\|_{\text{TV}}=\frac{1}{2}\sum_{x=1}^{d}|\mu(x)-\nu(x)|$
denotes the total variation distance and $\varepsilon>0$ is a prescribed tolerance.
Using the eigendecomposition $P=U^{-1}\Sigma U$, we can write
$\psi_{0}P^{t}(x)=\sum_{j=1}^{d}\bigl(\psi_{0}\cdot U^{-1}(\cdot,j)\bigr)\,\sigma_{j}^{t}\,U(j,x),$
where $\psi_{0}$ is treated as a row vector and $\psi_{0}\cdot U^{-1}(\cdot,j)$ denotes the scalar product with the $j$th column of $U^{-1}$. Using $\psi_{0}\cdot U^{-1}(\cdot,1)=1$ and $U(1,\cdot)=\pi$, we obtain
$\|\psi_{0}P^{t}-\pi\|_{\text{TV}}$ $=\frac{1}{2}\sum_{x=1}^{d}\left|\pi(x)-\sum_{j=1}^{d}\bigl(\psi_{0}\cdot U^{-1}(\cdot,j)\bigr)\,\sigma_{j}^{t}\,U(j,x)\right|$
$=\frac{1}{2}\sum_{x=1}^{d}\left|\pi(x)-\pi(x)+\sum_{j=2}^{d}\bigl(\psi_{0}\cdot U^{-1}(\cdot,j)\bigr)\,\sigma_{j}^{t}\,U(j,x)\right|$
$\leq\frac{1}{2}\sum_{x=1}^{d}\sum_{j=2}^{d}\bigl|\psi_{0}\cdot U^{-1}(\cdot,j)\bigr|\,|\sigma_{j}^{t}|\,|U(j,x)|$
$=\sum_{j=2}^{d}\sigma_{j}^{t}\,a_{j}\;\leq\;\sigma^{t}_{2}\sum_{j=2}^{d}a_{j},$
where $a_{j}\geq 0$ are constants that depend on $\psi_{0}$ and on the eigenvectors of $P$. Therefore, the total variation distance decays at least as fast as $\sigma_{2}^{t}$, and we obtain the upper bound
Since $0\leq\sigma_{2}<1$, the denominator $\log(\sigma_{2})$ is negative, and the right-hand side is positive. The smaller the value of $\sigma_{2}$, the faster the Markov chain mixes, and the more rapidly the autocorrelation $r_{yy}[k]$ decays.
In summary, by empirically estimating the long-lag decay of the autocorrelation function $r_{yy}[k]$ for some informative observable $f(x)$, we effectively probe the second-largest eigenvalue $\sigma_{2}$ of the transition kernel and thus obtain a proxy for the mixing time of the Gibbs sampler underlying our DTM.
Appendix H Total correlation penalty
@@
H.1 Gradients of the total correlation penalty
@@
H.2 Adaptive Correlation Penalty
@@
To address this, we employ an Adaptive Correlation Penalty (ACP) scheme that dynamically adjusts $\lambda_{t}$ based on an estimate of the model’s current mixing time. We use the autocorrelation of the Gibbs sampling chain, $r_{yy}^{t}$, as a proxy for mixing, as described in Section G and the main text, Eq. 18.
@@
Figure 14: The adaptive correlation penalty The dynamics of $r_{yy}^{t}$ and $\lambda_{t}$ over a training run. Large values of $r_{yy}^{t}$ lead to increasingly large values of $\lambda_{t}$, which cause $r_{yy}^{t}$ to decrease. The system reaches a stable configuration by the end of training.
Appendix I Embedding integers into Boltzmann machines
@@
Appendix J Deterministic embeddings for DTMs
@@
In section˜V of the paper, we mention hybrid thermodynamic models, the purpose of which is to combine the flexibility of conventional neural networks (NNs) with the efficiency of probabilistic computers. For example, in the context of image generation, a small convolutional neural network can be used to map color images into a format compatible with a binary DTM. To properly take advantage of the DTM’s energy efficiency, the conventional model should be at least 2 or 3 orders of magnitude smaller (e.g., in terms of parameter count or number of operations per sample) than the DTM.
@@
There are various options for the type of conventional model one can use for the embedding, e.g., invertible models such as GLOW [5] or Normalizing Flows [10], as well as simpler solutions, such as an Autoencoder.
@@
Our hybrid model achieved a FID score of $\sim 60$ on CIFAR10. Our DTM had 8 million parameters, the decoder had 65k, and the encoder and critic were both below 500k parameters. At inference time, only the DTM and the decoder are used. To achieve a similar performance with a conventional GAN, the decoder/generator requires about 500000 parameters.
Appendix K Some details on our RNG
@@
Appendix L MEBM experiments
Our experiments on MEBMs were conducted in the typical way [12]. We employed the same Boltzmann machine architecture as we typically use for the DTM layers, specifically $L=70$ with $G_{12}$ connectivity. Random nodes were chosen to represent the data, and the rest were left as latent variables (as discussed in section D).
@@
The exponential decay rates extracted from Fig. 16 were used as the mixing times in Fig. 2 in the article. Calling this a "mixing time" is a slight abuse of nomenclature. However, we did not think it made enough of a difference to the article’s message to disambiguate (since it is an upper bound on the mixing time, as discussed in Section G).
@@
Appendix M Sampling time vs Performance Tradeoff
The total time it takes to sample from a DTM is $O(KT)$, where $T$ is the number of denoising steps comprising the DTM and $K$. While one might use a different number of Gibbs steps at training time and inference time, it should be noted that for DTMs trained with adaptive correlation penalty (ACP), $K_{\text{inference}}$ should be similar or only slightly higher to $K_{\text{train}}$. This is substantially different from the usual good practice for un-penalised EBMs. The main reason for this distinction is ACP. In particular, when training an EBM (or DTM) with ACP, the closed-loop control used in ACP will make sure that the mixing time $K_{\text{mix}}$ of the EBM is approximately equal to $K_{\text{train}}$. It is always the case that using $K_{\text{inference}}>>K_{\text{mix}}$ brings no benefits compared to $K_{\text{inference}}\approx K_{\text{mix}}$ and therefore when sampling from an ACP-regularised model, using $K_{\text{inference}}\approx K_{\text{train}}$. On the other hand, for usual EBMs trained without ACP (or a similar penalty), the mixing time of the EBM can greatly exceed $K_{\text{train}}$. Due to the mixing-expressivity tradeoff, this at first means that the performance of this un-penalised EBM can exceed that of an ACP EBM, as long as $K_{\text{inference}}\geq K_{\text{mix}}>>K_{\text{train}}$. However, if we then keep training with $K_{\text{train}}<<K_{\text{mix}}$, the performance of the EBM starts to degrade substantially, as shown in fig.˜18. One may attempt to circumvent this by treating the EBM as a non-equilibrium model as explored in [Decelle_2022], but that introduces significant new complexity and is mathematically not well understood.
That being said, why does increasing $K_{\text{train}}$ improve the performance of an ACP-regularised DTM or MEBM? ACP works by periodically checking whether $K_{\text{mix}}>K_{\text{train}}$. If yes, then it penalizes the model more, reducing $K_{\text{mix}}$, but potentially hurting the expressivity of the model. If not, then it reduces the penalty coefficient, allowing both $K_{\text{mix}}$ and model expressivity to increase. Therefore, using a larger $K_{\text{train}}$ lets ACP be more lenient, resulting in a more expressive model.
Figure 17: The FID scores of different DTMs trained on binarised FashionMNIST. On the horizontal axis we vary the number of denoising steps $T$ and on the vertical axis we vary the number of Gibbs steps during training $K_{\text{train}}$. Each diagonal (running from top-left to bottom-right) represents a constant amount of energy expenditure. For this plot, the number of Gibbs steps at inference was twice the number used during training. The reported FID values were averaged over three runs with different random seeds.
Figure 18: FID and mixing time of an MEBM (same topology as most other experiments in this article) throughout training. As the MEBM trains and becomes more expressive (FID goes down) the mixing time increases. Around epoch 20 the model starts freezing (the mixing time becomes too long), which causes the CD gradients to become inaccurate and soon after the performance of the model starts to degrade. The mixing time subsequently also decreases somewhat, although this is likely an artefact of mode collapse rather than a sign of relaxation of the energy barriers in the model.
Appendix N CIFAR-10 Images
Figure 19: Images generated by the hybrid model trained on CIFAR-10 as described in section˜V and Appendix J. Images come in groups of 8, where each group shows the progression of the image throughout the denoising process. The denoising process operates inside a binary latent space, which is then converted into images by a small neural-network-based decoder.